Suppose the current USD/euro exchange rate is 1.2000 dollar per euro. The six month forward exchange rate is 1.1950. The six month USD interest rate is 1% per annum continuously compounded. Estimate the six month euro interest rate (expressed continuously compounded). Assume six months is 0.5 years.
Answer #1

Hi

Here six month euro interest rate =rf

Forward rate F = 1.1950

spot rate S = 1.

F = S*e^(rs-rf)*0.5

1.1950 = 1.2*e^(0.01-rf)*0.5

e^(0.01-rf) = (1.1950/1.2)^2

0.01-rf = ln(1.1950/1.2)^2

0.01-rf = 2ln (1.1950/1.2)

0.01 -rf = 2*(-0.004175)

0.01-rf = -0.008350

rf = 0.01835

rf = 1.835%

Hence six month euro interest rate is 1.835%

Thanks